Solve using the five-step problem-solving process. Show all steps necessary to arrive at your solution.
A mirror 2 feet by 6 feet is to be mounted on a door so that there is a uniform edge between the mirror and the door all around the mirror. How wide will the edge between the mirror and the door be if the area of the edge is 9 square feet?

Answer by checkley77(12844)

2*6=12 FT^2 IS THE AREA OF THE MIRROR.
(2+2X)(6+2X)=12+9
12+12X+4X+4X^2=21
4X^2+16X+12-21=0
4X^2+16X-9=0
(2X-1)(2X+9)=0
2X-1=0
2X=1
X=1/2 ANS.
2X+9=0
2X=-9
X=-9/2
X=-4.5 ANS.